# Best weighted approximation of some kernels on the real axis

### v1

@@ 1 Introduction
${\mathcal{K}}_{\lambda,s}(t):={\mathcal{K}}_{\lambda,s}(t,A,B)={\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}}},$ (1.1)
where $A$ and $B$ are fixed complex numbers, $\lambda>0$, and $s$ is a positive integer ($s\in\mathbb{N}$). A lot of works are devoted to the solutions of extremal problems of the best approximation of kernels of various types in uniform and integral metrics (see, for example, the monographs [1, 4, 13] and the paper [9], which provides a bibliography on this topic). In particular, Ahiezer [1] was the first to establish the exact values of the best weighted approximations on the real axis of kernels (1.1) in the case of $s=0$ and found the extreme polynomials. He proved that for any fixed algebraic polynomial
@@
Dzhrbashyan [3] developed a method that allowed obtaining solutions to extremal problems about the best rational approximations of the Cauchy kernel $\mathcal{C}:\mathbb{C}^{2}\rightarrow\mathbb{C}$
$(z,t)\mapsto{\textstyle\frac{1}{z-t}},\quad\mathrm{Im}~z\not=0,$
on the entire real axis $\infty<t<\infty$ in both uniform and mean square metrics. The proposed method was based on the use of certain orthogonal and bi-orthogonal systems of rational functions with fixed poles on the entire real axis $\mathbb{R}$.
Using the method of Dzhrabashyan, Voskanian in his paper [15] improved upon the result of Akhiezer, solved similar extremal problems for kernels (1.1) in the case of $s=0$ in the mean square metric, and also considered the corresponding problems of approximating such kernels using rational functions with fixed poles.
@@
For $A=s=0$ and $B={\textstyle\frac{1}{\pi}}$, the kernel ${\mathcal{K}}_{\lambda,s}(t,A,B)$ coincides with the Poisson kernel $P_{t}(\lambda)={\textstyle\frac{1}{\pi}}{\textstyle\frac{t}{\lambda^{2}+t^{2}}}$ as function of the variable $\lambda$.
@@
$\nabla^{2}(\nabla^{2}U)=0,\quad\nabla^{2}:={\textstyle\frac{\partial^{2}}{\partial x^{2}}}+{\textstyle\frac{\partial^{2}}{\partial\lambda^{2}}},$
@@
The approximation properties of similar Poisson integrals were investigated in [5, 12]. We also note that in the paper [9], weighted combinations of the Cauchy-Szegö kernel and its derivatives on the unit circle were studied, and, in turn, kernels (1.1) can be considered as a special case of linear combinations of Poisson kernels and their derivatives. Incidentally, the properties of Poisson kernels are often used in solving problems in physics, mechanics, electrical engineering, etc.
@@ 2 Preliminaries
For arbitrary $z,\zeta\in\mathbb{C},$ $z\not=\overline{\zeta}$, and any positive integer $n$
@@
$\nu_{k}(z,{\bf a}):=\sum_{|{\bf k}|=k}\prod_{j=1}^{n}{\textstyle\frac{1}{(z-a_{j})^{k_{j}}}},$ (2.3)
@@ 3 Main results
and by Leibnitz rule,
$\Big[{\textstyle\frac{1}{\tau_{n}(t,{\bf\bar{a}})}}\Big]^{(k)}_{t}=\sum_{|{\bf k}|=k}{\textstyle\frac{k!}{k_{1}!\cdot k_{2}!\cdot\ldots\cdot k_{n}!}}\prod_{j=1}^{n}\Big[{\textstyle\frac{1}{\tau_{n}(t,{\bf\bar{a}})}}\Big]^{(k_{j})}=\sum_{|{\bf k}|=k}{\textstyle\frac{k!}{k_{1}!\cdot k_{2}!\cdot\ldots\cdot k_{n}!}}\prod_{j=1}^{n}{\textstyle\frac{(-1)^{k_{j}}k_{j}!}{(t-\overline{a}_{j})^{k_{j}+1}}}={\textstyle\frac{(-1)^{k}k!\,\nu_{k}(t,{\bf\bar{a}})}{\tau_{n}(t,{\bf\bar{a}})}},$
@@
$\left[{\textstyle\frac{A+Bt}{(t+{\mathrm{i}}\lambda)^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}\right]^{(s)}_{t}$
$=$ $\sum_{|{\bf s}|=s}{\textstyle\frac{(A+Bt)s!}{s_{1}!\cdot s_{2}!\cdot s_{3}!}}{\textstyle\frac{(-1)^{s_{1}+1}{s_{1}}!}{(t-z)^{s_{1}+1}}}{\textstyle\frac{(-1)^{s_{3}}(s+s_{3})!}{s!(t+{\mathrm{i}}\lambda)^{s+s_{3}+1}}}{\textstyle\frac{(-1)^{s_{2}}{s_{2}}!\,\nu_{s_{2}}(t,{\bf\bar{a}})}{\tau_{n}(t,{\bf\bar{a}})}}$
$+\sum_{|{\bf s}|=s-1}\textstyle{{\textstyle\frac{B(s-1)!}{s_{1}!\cdot s_{2}!\cdot s_{3}!}}{\textstyle\frac{(-1)^{s_{1}+1}s_{1}!}{(t-z)^{s_{1}+1}}}{\textstyle\frac{(-1)^{s_{3}}(s+s_{3})!}{s!(t+{\mathrm{i}}\lambda)^{s+s_{3}+1}}}{\textstyle\frac{(-1)^{s_{2}}s_{2}!\,\nu_{s_{2}}(t,{\bf\bar{a}})}{\tau_{n}(t,{\bf\bar{a}})}}}$
$=$ ${\textstyle\frac{(-1)^{s+1}s!}{\tau_{n}(t,{\bf\bar{a}})}}\Bigg[\sum_{|{\bf s}|=s}{\textstyle\frac{(A+Bt){s+s_{3}\choose s}\nu_{s_{2}}(t,{\bf\bar{a}})}{(t-z)^{s_{1}+1}(t+{\mathrm{i}}\lambda)^{s+s_{3}+1}}}-\sum_{|{\bf s}|=s-1}{\textstyle\frac{B{s+s_{3}\choose s}\nu_{s_{2}}(t,{\bf\bar{a}})}{s(t-z)^{s_{1}+1}(t+{\mathrm{i}}\lambda)^{s+s_{3}+1}}}\Bigg],$
@@
$\left[{\textstyle\frac{A+Bt}{(t+{\mathrm{i}}\lambda)^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}\right]^{(s)}_{t}$ $=$ ${\textstyle\frac{(-1)^{s+1}s!}{\tau_{n}(t,{\bf\bar{a}})}}\Bigg[\sum_{s_{1}=0}^{s}\sum_{s_{2}=0}^{s-s_{1}}{\textstyle\frac{(A+Bt){2s-s_{1}-s_{2}\choose s}\nu_{s_{2}}(t,{\bf\bar{a}})}{(t-z)^{s_{1}+1}(t+{\mathrm{i}}\lambda)^{2s-s_{1}-s_{2}+1}}}$
@@
$\left[{\textstyle\frac{A+Bt}{(t+{\mathrm{i}}\lambda)^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}\right]^{(s)}_{t}$ $=$ ${\textstyle\frac{(-1)^{s+1}s!}{\tau_{n}(t,{\bf\bar{a}})}}\Bigg[\sum_{s_{1}=0}^{s}\sum_{s_{2}=0}^{s-s_{1}}{\textstyle\frac{(A+Bt){2s-s_{1}-s_{2}\choose s}\nu_{s_{2}}(t,{\bf\bar{a}})}{(t-z)^{s_{1}+1}(t+{\mathrm{i}}\lambda)^{2s-s_{1}-s_{2}+1}}}$
@@
$=$ ${\textstyle\frac{(-1)^{s+1}s!}{\tau_{n}(t,{\bf\bar{a}})}}\sum_{s_{1}=0}^{s}\sum_{s_{2}=0}^{s-s_{1}}{\textstyle\frac{{2s-s_{1}-s_{2}\choose s}\nu_{s_{2}}(t,{\bf\bar{a}})}{(t-z)^{s_{1}+1}(t+{\mathrm{i}}\lambda)^{2s-s_{1}-s_{2}+1}}}$
@@
$\left[{\textstyle\frac{A+Bt}{(t+{\mathrm{i}}\lambda)^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}\right]^{(s)}_{t}\!\!={\textstyle\frac{(-1)^{s+1}}{\tau_{n}(t,{\bf\bar{a}})}}\sum_{l=0}^{s}\sum_{j=0}^{s-l}{\textstyle\frac{s!\,{s+j\choose s}\,\nu_{s-l-j}(t,{\bf\bar{a}})}{(t-z)^{l+1}(t+{\mathrm{i}}\lambda)^{s+j+1}}}\Big(A+Bt-{\textstyle\frac{B(t+{\mathrm{i}}\lambda)j)}{s(s+j)}}\Big).$
@@
$\mathop{\rm res}\limits_{t=t_{2}}\textstyle{{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}}=\textstyle{{\textstyle\frac{(-1)^{s+1}}{\tau_{n}(t,{\bf\bar{a}})}}}\displaystyle{\sum_{l=0}^{s}\sum_{j=0}^{s-l}}\textstyle{{\textstyle\frac{{s+j\choose s}\nu_{s-l-j}({\mathrm{i}}\lambda,{\bf\bar{a}})}{({\mathrm{i}}\lambda-z)^{l+1}(2{\mathrm{i}}\lambda)^{s+j+1}}}\Big(A+B{\mathrm{i}}\lambda-{\textstyle\frac{2B{\mathrm{i}}\lambda j}{s(s+j)}}\Big)}.$
@@
$\mathop{\rm res}\limits_{t=t_{2}}{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}\!\!$ $=$ $\!\!\sum_{l=0}^{s}\sum_{j=0}^{s-l}{\textstyle\frac{{s+j\choose s}(2{\mathrm{i}}\lambda)^{s-j}\nu_{s-l-j}({\mathrm{i}}\lambda,{\bf\bar{a}})}{(2\lambda)^{2s+1}\tau_{n}({\mathrm{i}}\lambda,{\bf\bar{a}})({\mathrm{i}}\lambda-z)^{l+1}}}\Big({\textstyle\frac{B\lambda(2j-s(s+j))}{s(s+j)}}+{\mathrm{i}}A\Big)$ (3.7)
@@
$U_{n}^{\lambda}(z)={\textstyle\frac{1}{2\pi{\mathrm{i}}}}\int_{-\infty}^{\infty}{\textstyle\frac{\overline{B_{n+1}(t)}R_{n,s}^{\lambda}(t)}{t-z}}{\mathrm{d}}t={\textstyle\frac{\prod\limits_{k=1}^{n}\overline{\chi}_{k}}{2\pi{\rm i}}}\int_{-\infty}^{\infty}\overline{{\textstyle\frac{\tau_{n}(t,{\bf a})}{\tau_{n}(t,{\bf\bar{a}})}}}{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}\tau_{n}(t,{\bf\bar{a}})}}{\textstyle\frac{{\mathrm{d}}t}{t-z}}.$
@@
$\overline{{\textstyle\frac{\tau_{n}(t,{\bf a})}{\tau_{n}(t,{\bf\bar{a}})}}}={\textstyle\frac{\prod\limits_{k=1}^{n}\overline{(t-a_{k})}}{\prod\limits_{k=1}^{n}\overline{(t-\overline{a}_{k})}}}={\textstyle\frac{\prod\limits_{k=1}^{n}(t-\overline{a}_{k})}{\prod\limits_{k=1}^{n}(t-a_{k})}}={\textstyle\frac{\tau_{n}(t,{\bf\bar{a}})}{\tau_{n}(t,{\bf a})}}.$
@@
$\mathop{\rm res}\limits_{t=-{\mathrm{i}}\lambda}{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}(t-z)\tau_{n}(t,{\bf a})}}$
$=$ $\Bigg[\sum_{l=0}^{s}\sum_{j=0}^{s-l}{\textstyle\frac{{s+j\choose s}(-2{\mathrm{i}}\lambda)^{s-j}{\nu}_{s-l-j}(-{\mathrm{i}}\lambda,{\bf a})}{(2\lambda)^{2s+1}\tau_{n}(-{\mathrm{i}}\lambda,{\bf a})(-{\mathrm{i}}\lambda-z)^{l+1}}}\Big({\textstyle\frac{B\lambda(s(s+j)-2j)}{s(s+j)}}+{\mathrm{i}}A\Big)\Bigg]$
@@
${\mathcal{K}}_{\lambda,s}(z)-p_{n-1}^{(\lambda,s)}(z)=\sum_{l=0}^{s}\Big(\tau_{n}(z,{\bf\bar{a}}){\textstyle\frac{D_{l}}{({\mathrm{i}}\lambda-z)^{l+1}}}+\tau_{n}(z,{\bf a}){\textstyle\frac{\overline{D}_{l}}{(-{\mathrm{i}}\lambda-z)^{l+1}}}\Big),$ (3.12)
@@
$\mathcal{E}_{n}^{2}({\mathcal{K}}_{\lambda,s})_{2,\rho_{n}}={\textstyle\frac{4\pi}{(2\lambda)^{2s+3}\mu_{n}(\lambda,{\bf a})|\rho_{0}|^{2}}}\sum_{k=0}^{s}\sum_{l=0}^{s}\textstyle{{l+k\choose l}}G_{k}\overline{G}_{l},$ (3.14)
where $\mu_{n}(\lambda,{\bf a})=\prod\limits_{k=1}^{n}\Big[\alpha_{k}^{2}+(\beta_{k}+\lambda)^{2}\Big]$, $a_{j}=\alpha_{j}+{\mathrm{i}}\beta_{j}$, $\beta_{j}>0$, $j=1,2,\ldots,n$,
@@
is satisfied when $\lambda\geq 1$ or when $0<\lambda<1$ and the numbers $a_{k}$ belong to the set ${\mathbb{C}}_{+}\setminus\{z\in{\mathbb{C}}:|z|<1\}$.
@@
Similarly, for the product $(-{\mathrm{i}})^{k}\overline{\nu}_{k}({\mathrm{i}}\lambda,{\bf{a}})$ we have
$(-{\mathrm{i}})^{k}\overline{\nu}_{k}({\mathrm{i}}\lambda,{\bf{a}})=\sum_{|{\bf k}|=k}\prod_{j=1}^{n}z_{j}^{k_{j}}=\sum_{|{\bf k}|=k}{\bf r}^{\bf k}\Big(\cos({\bf k}\cdot\mbox{${\bf\varphi}$})+{\mathrm{i}}\sin({\bf k}\cdot\mbox{${\bf\varphi}$})\Big).$
@@
$\sum_{k=0}^{s}\sum_{l=0}^{s}\textstyle{{l+k\choose l}}G_{k}\overline{G}_{l}={\displaystyle\int_{0}^{\infty}}{\rm e}^{-t}{\displaystyle\sum_{k=0}^{s}\sum_{l=0}^{s}}{\textstyle\frac{G_{k}t^{k}}{k!}}{\textstyle\frac{\overline{G}_{l}t^{l}}{l!}}{\rm d}t={\displaystyle\int_{0}^{\infty}}{\rm e}^{-t}\Big|{\displaystyle\sum_{k=0}^{s}}{\textstyle\frac{G_{k}t^{k}}{k!}}\Big|^{2}{\rm d}t.$
@@
$\int_{-\infty}^{\infty}\left|{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}}}-p(t)\right|^{2}{\textstyle\frac{\mathrm{d}t}{|\rho_{n}(t)|^{2}}}={\textstyle\frac{1}{|\rho_{0}|^{2}}}\int_{-\infty}^{\infty}\Big|R_{n,s}^{\lambda}(z)-\sum_{k=1}^{n}b_{k}\Phi_{k}(t)\Big|^{2}\mathrm{d}t.$ (3.16)
Moreover, due to Parseval’s identity, for any polynomial $\sum\limits_{k=1}^{n}b_{k}\Phi_{k}(t)$ with respect to the system $\{\Phi_{k}(z)\}_{k=1}^{\infty}$, we have
$\int_{-\infty}^{\infty}\Big|R_{n,s}^{\lambda}(z)-\sum_{k=1}^{n}b_{k}\Phi_{k}(t)\Big|^{2}\mathrm{d}t\geq\int_{-\infty}^{\infty}|R_{n}^{\lambda}(t)-S_{n}(R_{n}^{\lambda})(t)|^{2}~\mathrm{d}t=:I$ (3.17)
@@
$\int_{-\infty}^{\infty}\left|{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}}}-p_{n-1}^{(\lambda,s)}(t)\right|^{2}{\textstyle\frac{\mathrm{d}t}{|\rho_{n}(t)|^{2}}}={\textstyle\frac{1}{|\rho_{0}|^{2}}}\int_{-\infty}^{\infty}\Big|R_{n,s}^{\lambda}(t)-S_{n}(R_{n}^{\lambda})(t)\Big|^{2}\mathrm{d}t={\textstyle\frac{I}{|\rho_{0}|^{2}}}.$

### v2

@@ 1 Introduction
${\mathcal{K}}_{\lambda,s}(t):={\mathcal{K}}_{\lambda,s}(t,A,B)={\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}}}$ (1.1)
where $A$ and $B$ are fixed complex numbers, $\lambda>0$, and $s$ is a positive integer ($s\in\mathbb{N}$). A lot of works are devoted to the solutions of extremal problems of the best approximation of kernels of various types in uniform and integral metrics (see, for example, [1, 4, 9, 13], which provide a bibliography on this topic). In particular, Ahiezer [1] was the first to establish the exact values of the best weighted approximations on the real axis of kernels (1.1) in the case of $s=0$ and found the extreme polynomials. He proved that for any fixed algebraic polynomial
@@
The Cauchy kernel is the function
$(z,t)\mapsto\frac{1}{z-t},\quad(z,t)\in\mathbb{C}^{2}\setminus\Delta,$
where $\Delta:=\{(z,t)\in\mathbb{C}^{2}:z=t\}$.
Dzhrbashyan [3] developed a method that provides solutions to extremal problems about the best rational approximations of the Cauchy kernel in the case where $z$ is fixed with $\mathrm{Im}z\not=0$ and $\infty<t<\infty$, both in the uniform and mean-square metrics on the real axis $\mathbb{R}$. The method is based on the use of certain orthogonal and bi-orthogonal systems of rational functions on $\mathbb{R}$ with fixed poles in the upper half-plane $\mathbb{C}_{+}:=\{z\in\mathbb{C}:\mathrm{Im}z>0\}$.
Using the method of Dzhrabashyan, Voskanian [15] improved upon the result of Akhiezer, solved similar extremal problems for kernels (1.1) in the case of $s=0$ in the mean square metric, and also considered the corresponding problems of approximating such kernels by rational functions with fixed poles.
@@
For $A=s=0$ and $B={\textstyle\frac{1}{\pi}}$, the kernel ${\mathcal{K}}_{\lambda,s}(t,A,B)$ coincides with the Poisson kernel $P_{t}(\lambda)={\textstyle\frac{1}{\pi}}{\textstyle\frac{t}{\lambda^{2}+t^{2}}}$ as a function of the variable $\lambda$.
@@
$\nabla^{2}(\nabla^{2}U)=0,\quad\nabla:={\textstyle\frac{\partial^{2}}{\partial x^{2}}}+{\textstyle\frac{\partial^{2}}{\partial\lambda^{2}}},$
@@
The approximation properties of similar Poisson integrals were investigated in [5, 12]. We also note that in [9], weighted combinations of the Cauchy-Szegö kernel and its derivatives on the unit circle were studied, and, in turn, kernels (1.1) can be considered as a special case of linear combinations of Poisson kernels and their derivatives. Incidentally, the properties of Poisson kernels are often used in solving problems in physics, mechanics, electrical engineering, etc.
@@ 2 Preliminaries
For arbitrary $z,\zeta\in\mathbb{C},$ $z\not=\overline{\zeta}$ and any positive integer $n$
@@
$\nu_{k}(z,{\bf a}):=\sum_{|{\bf k}|=k}\prod_{j=1}^{n}{\textstyle\frac{1}{(z-a_{j})^{k_{j}}}}.$ (2.3)
@@ 3 Main results
and by Leibniz rule,
$\Big[{\textstyle\frac{1}{\tau_{n}(t,{\bf\bar{a}})}}\Big]^{(k)}_{t}=\sum_{|{\bf k}|=k}{\textstyle\frac{k!}{k_{1}!k_{2}!\cdots k_{n}!}}\prod_{j=1}^{n}\Big[{\textstyle\frac{1}{t-\overline{a}_{j}}}\Big]^{(k_{j})}=\sum_{|{\bf k}|=k}{\textstyle\frac{k!}{k_{1}!k_{2}!\cdots k_{n}!}}\prod_{j=1}^{n}{\textstyle\frac{(-1)^{k_{j}}k_{j}!}{(t-\overline{a}_{j})^{k_{j}+1}}}={\textstyle\frac{(-1)^{k}k!\,\nu_{k}(t,{\bf\bar{a}})}{\tau_{n}(t,{\bf\bar{a}})}},$
@@
$\left[{\textstyle\frac{A+Bt}{(t+{\mathrm{i}}\lambda)^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}\right]^{(s)}_{t}$ $=$ $\sum_{|{\bf s}|=s}{\textstyle\frac{(A+Bt)s!}{s_{1}!s_{2}!s_{3}!}}{\textstyle\frac{(-1)^{s_{1}}{s_{1}}!}{(t-z)^{s_{1}+1}}}{\textstyle\frac{(-1)^{s_{3}}(s+s_{3})!}{s!(t+{\mathrm{i}}\lambda)^{s+s_{3}+1}}}{\textstyle\frac{(-1)^{s_{2}}{s_{2}}!\,\nu_{s_{2}}(t,{\bf\bar{a}})}{\tau_{n}(t,{\bf\bar{a}})}}$
$+\sum_{|{\bf s}|=s-1}\textstyle{{\textstyle\frac{B(s-1)!}{s_{1}!s_{2}!s_{3}!}}{\textstyle\frac{(-1)^{s_{1}}s_{1}!}{(t-z)^{s_{1}+1}}}{\textstyle\frac{(-1)^{s_{3}}(s+s_{3})!}{s!(t+{\mathrm{i}}\lambda)^{s+s_{3}+1}}}{\textstyle\frac{(-1)^{s_{2}}s_{2}!\,\nu_{s_{2}}(t,{\bf\bar{a}})}{\tau_{n}(t,{\bf\bar{a}})}}}$
$=$ ${\textstyle\frac{(-1)^{s}s!}{\tau_{n}(t,{\bf\bar{a}})}}\Bigg[\sum_{|{\bf s}|=s}{\textstyle\frac{(A+Bt){s+s_{3}\choose s}\nu_{s_{2}}(t,{\bf\bar{a}})}{(t-z)^{s_{1}+1}(t+{\mathrm{i}}\lambda)^{s+s_{3}+1}}}-\sum_{|{\bf s}|=s-1}{\textstyle\frac{B{s+s_{3}\choose s}\nu_{s_{2}}(t,{\bf\bar{a}})}{s(t-z)^{s_{1}+1}(t+{\mathrm{i}}\lambda)^{s+s_{3}+1}}}\Bigg],$
@@
$\left[{\textstyle\frac{A+Bt}{(t+{\mathrm{i}}\lambda)^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}\right]^{(s)}_{t}$ $=$ ${\textstyle\frac{(-1)^{s}s!}{\tau_{n}(t,{\bf\bar{a}})}}\Bigg[\sum_{s_{1}=0}^{s}\sum_{s_{2}=0}^{s-s_{1}}{\textstyle\frac{(A+Bt){2s-s_{1}-s_{2}\choose s}\nu_{s_{2}}(t,{\bf\bar{a}})}{(t-z)^{s_{1}+1}(t+{\mathrm{i}}\lambda)^{2s-s_{1}-s_{2}+1}}}$
@@
$\left[{\textstyle\frac{A+Bt}{(t+{\mathrm{i}}\lambda)^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}\right]^{(s)}_{t}$ $=$ ${\textstyle\frac{(-1)^{s}s!}{\tau_{n}(t,{\bf\bar{a}})}}\Bigg[\sum_{s_{1}=0}^{s}\sum_{s_{2}=0}^{s-s_{1}}{\textstyle\frac{(A+Bt){2s-s_{1}-s_{2}\choose s}\nu_{s_{2}}(t,{\bf\bar{a}})}{(t-z)^{s_{1}+1}(t+{\mathrm{i}}\lambda)^{2s-s_{1}-s_{2}+1}}}$
@@
$=$ ${\textstyle\frac{(-1)^{s}s!}{\tau_{n}(t,{\bf\bar{a}})}}\sum_{s_{1}=0}^{s}\sum_{s_{2}=0}^{s-s_{1}}{\textstyle\frac{{2s-s_{1}-s_{2}\choose s}\nu_{s_{2}}(t,{\bf\bar{a}})}{(t-z)^{s_{1}+1}(t+{\mathrm{i}}\lambda)^{2s-s_{1}-s_{2}+1}}}$
@@
$\left[{\textstyle\frac{A+Bt}{(t+{\mathrm{i}}\lambda)^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}\right]^{(s)}_{t}\!\!={\textstyle\frac{(-1)^{s}}{\tau_{n}(t,{\bf\bar{a}})}}\sum_{l=0}^{s}\sum_{j=0}^{s-l}{\textstyle\frac{s!\,{s+j\choose s}\,\nu_{s-l-j}(t,{\bf\bar{a}})}{(t-z)^{l+1}(t+{\mathrm{i}}\lambda)^{s+j+1}}}\Big(A+Bt-{\textstyle\frac{B(t+{\mathrm{i}}\lambda)j}{s(s+j)}}\Big).$
@@
$\mathop{\rm res}\limits_{t=t_{2}}\textstyle{{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}}=\textstyle{{\textstyle\frac{(-1)^{s}}{\tau_{n}({\mathrm{i}}\lambda,{\bf\bar{a}})}}}\displaystyle{\sum_{l=0}^{s}\sum_{j=0}^{s-l}}\textstyle{{\textstyle\frac{{s+j\choose s}\nu_{s-l-j}({\mathrm{i}}\lambda,{\bf\bar{a}})}{({\mathrm{i}}\lambda-z)^{l+1}(2{\mathrm{i}}\lambda)^{s+j+1}}}\Big(A+B{\mathrm{i}}\lambda-{\textstyle\frac{2B{\mathrm{i}}\lambda j}{s(s+j)}}\Big)}.$
@@
$\mathop{\rm res}\limits_{t=t_{2}}{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}(t-z)\tau_{n}(t,{\bf\bar{a}})}}\!\!$ $=$ $\!\!\sum_{l=0}^{s}\sum_{j=0}^{s-l}{\textstyle\frac{{s+j\choose s}(2{\mathrm{i}}\lambda)^{s-j}\nu_{s-l-j}({\mathrm{i}}\lambda,{\bf\bar{a}})}{(2\lambda)^{2s+1}\tau_{n}({\mathrm{i}}\lambda,{\bf\bar{a}})({\mathrm{i}}\lambda-z)^{l+1}}}\Big({\textstyle\frac{B\lambda(s(s+j)-2j)}{s(s+j)}}-{\mathrm{i}}A\Big)$ (3.7)
@@
$U_{n}^{\lambda}(z)={\textstyle\frac{1}{2\pi{\mathrm{i}}}}\int_{-\infty}^{\infty}{\textstyle\frac{\overline{B_{n+1}(t)}R_{n,s}^{\lambda}(t)}{t-z}}{\mathrm{d}}t={\textstyle\frac{\prod\limits_{k=1}^{n}\overline{\chi}_{k}}{2\pi{\rm i}}}\int_{-\infty}^{\infty}\overline{\left({\textstyle\frac{\tau_{n}(t,{\bf a})}{\tau_{n}(t,{\bf\bar{a}})}}\right)}{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}\tau_{n}(t,{\bf\bar{a}})}}{\textstyle\frac{{\mathrm{d}}t}{t-z}}.$
@@
$\overline{\Big({\textstyle\frac{\tau_{n}(t,{\bf a})}{\tau_{n}(t,{\bf\bar{a}})}}\Big)}={\textstyle\frac{\prod\limits_{k=1}^{n}\overline{(t-a_{k})}}{\prod\limits_{k=1}^{n}\overline{(t-\overline{a}_{k})}}}={\textstyle\frac{\prod\limits_{k=1}^{n}(t-\overline{a}_{k})}{\prod\limits_{k=1}^{n}(t-a_{k})}}={\textstyle\frac{\tau_{n}(t,{\bf\bar{a}})}{\tau_{n}(t,{\bf a})}}.$
@@
$\mathop{\rm res}\limits_{t=-{\mathrm{i}}\lambda}{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}(t-z)\tau_{n}(t,{\bf a})}}$ $=$ $\Bigg[\sum_{l=0}^{s}\sum_{j=0}^{s-l}{\textstyle\frac{{s+j\choose s}(-2{\mathrm{i}}\lambda)^{s-j}{\nu}_{s-l-j}(-{\mathrm{i}}\lambda,{\bf a})}{(2\lambda)^{2s+1}\tau_{n}(-{\mathrm{i}}\lambda,{\bf a})(-{\mathrm{i}}\lambda-z)^{l+1}}}\Big({\textstyle\frac{B\lambda(s(s+j)-2j)}{s(s+j)}}+{\mathrm{i}}A\Big)\Bigg]$
@@
${\textstyle\frac{A+Bz}{(z^{2}+\lambda^{2})^{s+1}}}-p_{n-1}^{(\lambda,s)}(z)=\sum_{l=0}^{s}\Big(\tau_{n}(z,{\bf\bar{a}}){\textstyle\frac{D_{l}}{({\mathrm{i}}\lambda-z)^{l+1}}}+\tau_{n}(z,{\bf a}){\textstyle\frac{\overline{D}_{l}}{(-{\mathrm{i}}\lambda-z)^{l+1}}}\Big),$ (3.12)
@@
$\mathcal{E}_{n}^{2}({\mathcal{K}}_{\lambda,s})_{2,\rho_{n}}={\textstyle\frac{4\pi}{(2\lambda)^{2s+3}|\rho_{n}(-\mathrm{i}\lambda)|^{2}}}\sum_{k=0}^{s}\sum_{l=0}^{s}\textstyle{{l+k\choose l}}G_{k}\overline{G}_{l},$ (3.14)
where
@@
is satisfied when $\lambda>1$ or when $0<\lambda<1$ and the numbers $a_{k}$ belong to the set ${\mathbb{C}}_{+}\setminus\{z\in{\mathbb{C}}:|z|<1\}$.
@@
Similarly, for the product $(-{\mathrm{i}})^{k}\overline{\nu}_{k}({\mathrm{i}}\lambda,\bar{\bf{a}})$ we have
$(-{\mathrm{i}})^{k}\overline{\nu}_{k}({\mathrm{i}}\lambda,\bar{\bf{a}})=\sum_{|{\bf k}|=k}\prod_{j=1}^{n}z_{j}^{k_{j}}=\sum_{|{\bf k}|=k}{\bf r}^{\bf k}\Big(\cos({\bf k}\cdot\mbox{${\bf\varphi}$})+{\mathrm{i}}\sin({\bf k}\cdot\mbox{${\bf\varphi}$})\Big).$
@@
$\sum_{k=0}^{s}\sum_{l=0}^{s}\textstyle{{l+k\choose l}}G_{k}\overline{G}_{l}={\displaystyle\int_{0}^{\infty}}{\rm e}^{-t}{\displaystyle\sum_{k=0}^{s}\sum_{l=0}^{s}}{\textstyle\frac{G_{k}t^{k}}{k!}}{\textstyle\frac{\overline{G}_{l}t^{l}}{l!}}{\rm d}t={\displaystyle\int_{0}^{\infty}}{\rm e}^{-t}\Big|{\displaystyle\sum_{k=0}^{s}}{\textstyle\frac{G_{k}t^{k}}{k!}}\Big|^{2}{\rm d}t\geq 0.$
@@
$\int_{-\infty}^{\infty}\left|{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}}}-p(t)\right|^{2}{\textstyle\frac{\mathrm{d}t}{|\rho_{n}(t)|^{2}}}={\textstyle\frac{1}{|\rho_{0}|^{2}}}\int_{-\infty}^{\infty}\Big|R_{n,s}^{\lambda}(t)-\sum_{k=1}^{n}b_{k}\Phi_{k}(t)\Big|^{2}\mathrm{d}t.$ (3.16)
Moreover, due to the well-known minimum property of the partial sums of Fourier series, for any quasi-polynomial $\sum\limits_{k=1}^{n}b_{k}\Phi_{k}(t)$ with respect to the system $\{\Phi_{k}\}_{k=1}^{\infty}$, we have
$\int_{-\infty}^{\infty}\Big|R_{n,s}^{\lambda}(t)-\sum_{k=1}^{n}b_{k}\Phi_{k}(t)\Big|^{2}\mathrm{d}t\geq\int_{-\infty}^{\infty}|R_{n,s}^{\lambda}(t)-S_{n}(R_{n,s}^{\lambda})(t)|^{2}~\mathrm{d}t=:I$ (3.17)
@@
$\int_{-\infty}^{\infty}\left|{\textstyle\frac{A+Bt}{(t^{2}+\lambda^{2})^{s+1}}}-p_{n-1}^{(\lambda,s)}(t)\right|^{2}{\textstyle\frac{\mathrm{d}t}{|\rho_{n}(t)|^{2}}}={\textstyle\frac{1}{|\rho_{0}|^{2}}}\int_{-\infty}^{\infty}\Big|R_{n,s}^{\lambda}(t)-S_{n}(R_{n,s}^{\lambda})(t)\Big|^{2}\mathrm{d}t={\textstyle\frac{I}{|\rho_{0}|^{2}}}.$